OpenStudy (anonymous):

Order the numbers 0.64, 2/3 , 65%, and 7/10 from least to greatest.

OpenStudy (anonymous):

@CandyCove

candycove (candycove):

What do you think?

OpenStudy (anonymous):

2/3 0.64 65% 7/10

candycove (candycove):

We are supposed to covert the numbers in decimal form. 2/3 = 0.6666 65% = 65/100 = 0.65 7/10 = 0.7 Order: 0.64, 65%, 2/3, 7/10 ✔